OpenStudy (anonymous):

factor 4cos^2x-4sinx-5=0

OpenStudy (mertsj):

\[4\cos ^2x-4\sin x-5=0\] \[4(1-\sin ^2x)-4\sin x-5=0\] \[4-4\sin ^2x-4\sin x-5=0\]

OpenStudy (mertsj):

\[-4\sin ^2x-4sinx-1=0\]

OpenStudy (mertsj):

\[4\sin ^2x+4\sin x+1=0\]

OpenStudy (mertsj):

Factor that.

OpenStudy (anonymous):

how did you get a sin from a cos?